# A mechanism-informed reinforcement learning framework for shape optimization of airfoils

### v1

@@ Abstract
In this study, we present the mechanism-informed reinforcement learning framework for airfoil shape optimization. By leveraging the twin delayed deep deterministic policy gradient algorithm for its notable stability, our approach addresses the complexities of optimizing shapes governed by fluid dynamics. The PDEs-based solver is adopted for its accuracy even when the configurations and geometries are extraordinarily changed during the exploration. Dual-weighted residual-based mesh refinement strategy is applied to ensure the accurate calculation of target functionals. To streamline the iterative optimization process and handle geometric deformations, our approach integrates Laplacian smoothing, adaptive refinement, and a Bézier fitting strategy. This combination not only remits mesh tangling but also guarantees a precise manipulation of the airfoil geometry. Our neural network architecture leverages Bézier curves for efficient dimensionality reduction, thereby enhancing the learning process and ensuring the geometric accuracy of the airfoil shapes. An attention mechanism is embedded within the network to calculate potential action on the state as well. Furthermore, we have introduced different reward and penalty mechanisms tailored to the specific challenges of airfoil optimization. This algorithm is designed to support the optimization task, facilitating a more targeted and effective approach for airfoil shape optimization.
@@ 1 Introduction
Aerodynamic design has been an important research theme for decades, particularly in the context of optimizing airfoils for enhanced aerodynamic performance[1, 2] and fuel efficiency[3]. Multidisciplinary analysis and optimization plays an indispensable part in these fields. Nonetheless, it is pointed out by an important proposal from NASA [4] that high-fidelity computational fluid dynamics(CFD) is not routinely used in this area. Both academia and industry anticipate significant advancements in the integration of CFD methodologies by the year 2030. Aiming at conducting shape optimization of airfoils based on a high-fidelity CFD solver with the state-of-art technique, a mechanism-informed reinforcement learning framework is proposed in this paper.
@@
Classical methods, such as gradient-based techniques[5], genetic algorithms[6] can be adopted to support the optimization task. The traditional method pioneered by Jameson [7, 8] combined the optimization problem with adjoint equations which has been proven to be very efficient. However, the analysis and application of optimization techniques in the industry can not be generalized to more complicated target functional easily[4]. Combining CFD methods with optimization offers some distinct advantages. For instance, a more precise depiction of flow fields enables a more accurate calculation of target functionals. The advent of techniques such as the dual-weighted residual(DWR)-based mesh adaptation[9, 10, 11] allows for the attainment of more precise target functionals while preserving computational accuracy. Moreover, the enhanced coupling between more accurate geometric representations and PDEs provides superior tools for simulating airfoil geometric deformations through the use of parametric curves and unstructured mesh methods to describe shapes[12, 13]. However, PDEs-based optimization also presents various challenges. These include the high computational cost associated with solving PDEs[14] with complex geometries, non-linearities inherent in many physical systems, and application issues for ensuring that optimized shapes are physically realizable[15]. Despite these obstacles, the benefits of PDEs-based approaches such as their precision and robustness, make the technique a powerful tool in the shape optimization domain.
@@
With the development of artificial intelligence in recent years, associating the neural networks with traditional computational fluid dynamics problems becomes feasible in shape optimization of airfoil[16]. In recent years, researchers have been exploring the integration of machine learning into the domain of shape optimization. For example, the Bézier-GAN developed in [17, 18], has been utilized to parameterize aerodynamic designs by learning from shape variations. Similarly, the conditional variational autoencoder [19] has been employed to dissect the intricate relationship between the aerodynamic performance and the shape of airfoils. Moreover, the recent advancements in reinforcement methodologies, such as deep reinforcement learning[20], have significantly expanded its applicability and effectiveness in handling high-dimensional state and action spaces, which provide a basis for the PDEs-based optimization problems.
@@
The distinctive advantages of reinforcement learning, compared with other machine learning algorithms, lie in its ability to make sequential decisions and learn optimal policies directly from interaction with the environment[21]. This is especially pertinent in the realm of control problems governed by partial differential equations[22, 23], where the dynamics are complex and highly nonlinear. Unlike traditional machine learning approaches that often require labeled data sets for training, reinforcement learning thrives in an exploration-based learning environment, making it uniquely suited for optimizing shapes and controls in PDEs-driven systems without the need for extensive pre-processed data. These developments, combined with our group’s work in exploring advanced PDE solvers[24, 25, 26, 12], have laid the groundwork for implementing a PDE solver-based reinforcement learning framework.
@@
Even though reinforcement learning has made significant strides in various applications, its integration into shape optimization, particularly within the context of PDEs-based systems, remains nascent. Ever since in [27], reinforcement learning has been applied to morphing within an airfoil design context. Initially, the application was constrained by the consideration of only four variables for airfoil morphing, primarily due to the extensive time demands of the learning process. With the advent of computational advancements, a direct shape optimization methodology utilizing reinforcement learning was introduced in [28]. However, the number of parameters is still far from the application. Given that computational fluid dynamics solver operations significantly consume CPU time, [29] implemented a transfer learning technique to expedite the learning process. In our research[30, 31, 32], we have integrated the DWR-based mesh refinement strategy, a method that can generate a high quality mesh for the calculation of target functional. With the corresponding parallelization module constructed to accelerate the calculation process[33], our PDE solver can deal with the reinforcement learning process where thousands of airfoil shapes should be taken into consideration.
However, embarking on a shape optimization methodology based on a precise PDE solver, particularly within an unstructured mesh presents additional challenges as follows. i). Automating the iterative process of the entire algorithm, which is highlighted in [34, 35]. During the iterative process of reinforcement learning, the PDE solver might be invoked thousands of times. Manual intervention at each iteration is impractical. However, an automated strategy necessitates a profound understanding of the parameters and robustness throughout the iterative process. This means that even with significant changes in configuration and geometry, the simulation should remain effective. ii). Managing geometric deformations on unstructured meshes, which encompasses potential mesh tangling issues[36]. Mesh deformation is significant for the shape optimization problem, especially for the design of airfoil where the physical properties of the fluid medium are significantly influenced by the geometry of the computational domain. The process of mesh deformation entails the manipulation of mesh nodes and elements to accurately represent the geometry and to facilitate the resolution of complex physical phenomena with high precision. Even though meshless techniques[37] can be adopted, mesh-based approaches still hold a predominant position due to their rigorous theoretical analysis and more advanced method for calculating target functional[38]. However, the deformation on the unstructured meshes is not straightforward because the elements are sensitive to the movement of the boundary that depicts the airfoil. The change of geometry may be stochastic initially, where the quality of meshes can not be guaranteed. Besides, the curvature may change rapidly along the deformed shape. If the curvature variation occurs within the elements, it is challenging to capture the geometry correctly. iii). Tackling the challenges posed by controlling geometries with a large number of parameters, including their geometric representation and the control of high-dimensional parameters for reinforcement learning[39, 40]. The primary difficulty with a high parameter count is that it complicates the simulation process. Furthermore, while a more precise control over geometry enhances manipulability, it also poses challenges to maintaining a cohesive global shape. Excessive control points can lead to issues with the smoothness of curves, which in turn, impacts the reinforcement learning algorithms. For instance, even if the deformation trend is correct, the lack of geometric smoothness can negatively influence the agent’s strategy, potentially leading to the failure of policy iteration. iv). Developing a framework capable of addressing different optimization objectives. The effectiveness of DWR-based mesh adaptation methods is inherently related to the current state of theoretical advancements[9]. The challenge is further compounded when dealing with more complex target functionals, whose resolution poses significant difficulties. Furthermore, translating traditional optimization problems into reinforcement learning frameworks introduces additional complexities. Ensuring the equivalence of primal optimization objectives and the result of reinforcement learning is paramount. This requires the understanding of the reward function[41], necessitating it to be closely compatible with the underlying optimization context.
In this paper, we proposed a novel framework, the mechanism-informed reinforcement learning, for conducting the shape optimization design of airfoils. The problems mentioned above are addressed with techniques as follows. Firstly, in addressing the challenges of automation, the principal difficulty in coupling our framework with current DWR-based mesh adaptation PDE solvers lies in the automatic selection of appropriate tolerance values. These values guide the mesh refinement process across different geometries. In [33], we introduced an algorithm that is capable of dealing with such issues. It not only facilitates automation which is crucial for iterative optimization, but also conserves time while maintaining accuracy. Implementation has been realized within AFVM4CFD, a C++ library maintained by our group. Subsequently, for the deformation of airfoils, we employed Python scripts to interface with the underlying library, facilitating the collection of critical reinforcement learning metrics such as states, actions, and rewards.
Secondly, to address potential mesh tangling issues during the airfoil deformation process, we employed a Laplacian smoothing technique to restore mesh quality after each deformation. Typically, once mesh entanglement has occurred, it is not feasible to rectify the mesh to its optimal state using such methods. On one hand, the restoration process after each deformation helps prevent the accumulation of irregular triangles, significantly reducing the possibility of mesh tangling issues. On the other hand, from the perspective of finite element analysis, irregular triangles of unstructured meshes can introduce additional errors. To manipulate the geometry with more elaborate actions, we utilized over 200 sample points to represent the shape. However, for tensors denoting the shape with such dimensions, local point movements represent a minor modification to the tensors, making it challenging for algorithms to discern significant differences before and after deformation. We introduced a smoothing movement process, ensuring that sample points near the moving point adjust according to this trend. It amplifies the difference between tensors. Additionally, this process also plays a non-negligible role in mitigating mesh tangling. When mesh deformation occurs within the initial mesh simulation dimensions, we employ an adaptive refinement process to capture those elements with a larger geometric change. With such an adaptive refinement method, we are able to capture subtle curvature changes smaller than the mesh size.
Thirdly, to avoid the jagged structures caused by numerous minor curvature changes, we added an additional regularization correction to the process of generating control points from sample points. The control points are restricted to have as equidistant intervals as possible, which significantly makes the shape smoother. Furthermore, the process of generating control points through Bezier fitting serves as an efficient means of dimensionality reduction, allowing for finer geometric control. This also aids the stability of the learning process by facilitating dimensionality reduction.
Last but not least, in order to develop a framework suitable for tackling control problems of this nature, identifying an appropriate reinforcement learning algorithm is essential. We chose the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm for stability on continuous optimization problems. The algorithm should be carefully designed as shape optimal design problem can be highly non-linear and sensitive to action parameters. In terms of neural network design, the initial step involves the comprehensive extraction of airfoil shape features. To this end, we employed various techniques, including feature enhancement and leveraging the intrinsic physical properties of geometric points, such as thickness and curvature, to bolster understanding of geometry. Although we experimented with using Variational Autoencoders (VAEs) for feature extraction, the challenges associated with learning an appropriate shape in a high-dimensional space became evident. Consequently, utilizing Bézier curves to reduce dimensionality to a finite set of control points emerged as a more effective representation method. We incorporated specific mapping functions into the action parameters output by the actor network, acknowledging the geometric significance of actual action variations. However, this approach risked potential gradient vanishing issues. Therefore, the actor network was designed not to be overly deep, further underscoring the importance of dimensionality reduction. Techniques such as residual connections[42], the SELU activation function[43], and real-time monitoring via TensorBoard[44] were employed to mitigate and monitor potential gradient vanishing problems. The design of the critic network had to account for the potential impact of actions and their parameters on the high-dimensional state. Hence, the attention mechanism proved to be an effective tool for achieving this goal. For gradient updates, we sampled from the replay buffer, selecting the best historical data, those closest to the current moment, and a weighted random assortment to facilitate policy updates. This approach aids in navigating the optimization landscape more effectively by emphasizing the exploration of promising regions while avoiding local optima. The reward function design emanates from the original optimization problem, aligned with the existing learning framework to define an appropriate reward structure. In the execution phase, we focused on encouraging behaviors associated with higher rewards and penalizing actions that violate constraints, alongside employing empirical exploration strategies. These measures ensure a more efficient iterative process of the algorithm, guiding the learning towards fulfilling the optimization objectives effectively.
@@
The rest of this paper is organized as follows. In Section 2, we briefly introduced the fundamental notations and provided a concise introduction to the Newton-GMG solver. In Section 3, the architecture of our automatic DWR-based mesh adaptation method is introduced. In Section 4, the challenge of mesh deformation within the unstructured mesh is discussed and the recovering method is developed. In section 5, the algorithm of the reinforcement learning, Twin Delayed DDPG is constructed. In Section 6, we conduct the optimization task and present the numerical results.
2 A breif review of Newton-GMG framework
@@
To embark on research involving shape optimization through reinforcement learning, it’s imperative to have a suitable set of governing equations at our disposal. These equations help us understand the distribution of specific physical quantities across space, thereby enabling the precise calculation of the objective functional targeted for optimization. In our previous works [31, 30, 25], we have developed a solver capable of handling the steady Euler equations. This solver not only facilitates the computation of these equations but also ensures the accurate resolution of the target functional, laying a solid foundation for our shape optimization endeavors. A brief introduction will be given in this section.
2.1 Basic notations and finite volume discretization
@@
The steady Euler equations form the backbone of our understanding of inviscid flow behavior and are capable of dealing with the shape optimal design problem[45]. These equations, represented in a conservative form for the inviscid two-dimensional steady state,
@@
To address these equations numerically, we turn our focus to discretizing the domain $\Omega$. Consider $\Omega$ as a domain in $\mathbb{R}^{2}$ with boundary denoted by $\Gamma$. We divide $\Omega$ into distinct control volumes, $K_{i}$, via a shape-regular subdivision. Then the mesh under the current mesh size is denoted as $\mathcal{K}_{h}$. Each control volume is referred to as $K_{i}$. The intersection between any two elements $K_{i}$ and $K_{j}$, is labeled as $e_{i,j}$, satisfying $e_{i,j}=\partial K_{i}\cap\partial K_{j}$. The edge $e_{i,j}$ relative to $K_{i}$ is equipped with a unit outer normal vector, $n_{i,j}$.
This setup allows us to redefine the weak form of Euler equations for the discretized domain through the divergence theorem as:
@@
2.2 Newton-GMG solver
To tackle the nonlinear Equation (5), we employ the Newton method for linearization, supplemented by a linear multigrid method for solving the equations as proposed in [46]. Initially, we expand Equation (5) using the Taylor series and neglect higher-order terms, simplifying the equations to:
[…]
where $\frac{\partial\mathcal{H}(\cdot,\cdot,\cdot)}{\partial\mathbf{u}}$ represents the Jacobian matrix of the numerical flux, and $\Delta\mathbf{u}_{i}$ represents the increment of conservative variables for the $i$-th element. Following each Newton iteration, the cell average undergoes an update, $\mathbf{u}_{i}^{(n+1)}=\mathbf{u}_{i}^{(n)}+\Delta\mathbf{u}_{i}^{(n)}$.
A regularization term is incorporated into the simulation to enhance system stability. Initially, the solution substantially deviates from a steady state, but as iterations advance, it converges towards equilibrium, reducing the regularization term to near-zero.
Refer to captionRefer to caption
Figure 1: Mesh around the NACA0012 airfoil.
For solving Equation (6), the geometric multigrid technique is adopted. As demonstrated in [32], applying regularization and geometric multigrid methods to both primal and dual equations significantly enhances solver robustness, ensuring reliable dual equation resolution.
Refer to captionRefer to caption
3 Automatic DWR-based adaptation for the target functional
In our work as discussed in [30, 32], we elaborated on a framework innovated by Darmofal [47], which showcases a fully discrete adjoint-weighted residual technique. This approach was further extended to address two-dimensional inviscid flow scenarios as depicted in [48]. To incorporate this technique within our framework, we present an overview of the methodology.
We denote solutions on the coarse mesh as $\mathbf{u}_{H}$ and on the finer mesh as $\mathbf{u}_{h}$. The initial goal was to precisely compute the numerical integration of $J(\mathbf{u})$ over a fine mesh, denoted as $J_{h}(\mathbf{u}_{h})$ which needs to be solved on the fine mesh directly. Subsequently, to develop the extrapolation method that can calculate this value from the coarse mesh, a multi-variable Taylor series expansion method is adopted as follows:

### v2

# A mechanism-driven reinforcement learning framework for shape optimization of airfoils

@@ Abstract
In this paper, a novel mechanism-driven reinforcement learning framework is proposed for airfoil shape optimization. To validate the framework, a reward function is designed and analyzed, from which the equivalence between the maximizing the cumulative reward and achieving the optimization objectives is guaranteed theoretically. To establish a quality exploration, and to obtain an accurate reward from the environment, an efficient solver for steady Euler equations is employed in the reinforcement learning method. The solver utilizes the Bézier curve to describe the shape of the airfoil, and a Newton-geometric multigrid method for the solution. In particular, a dual-weighted residual-based h-adaptive method is used for efficient calculation of target functional. To effectively streamline the airfoil shape during the deformation process, we introduce the Laplacian smoothing, and propose a Bézier fitting strategy, which not only remits mesh tangling but also guarantees a precise manipulation of the geometry. In addition, a neural network architecture is designed based on an attention mechanism to make the learning process more sensitive to the minor change of the airfoil geometry. Numerical experiments demonstrate that our framework can handle the optimization problem with hundreds of design variables. It is worth mentioning that, prior to this work, there are limited works combining such high-fidelity partial differential equatons framework with advanced reinforcement learning algorithms for design problems with such high dimensionality.
@@ 1 Introduction
Aerodynamic design has been an important research theme for decades, particularly in the context of optimizing airfoils for enhanced aerodynamic performance[1, 2] and fuel efficiency[3]. Multidisciplinary analysis and optimization play an indispensable part in these fields. Nonetheless, it is pointed out by an important proposal from NASA [4] that high-fidelity computational fluid dynamics(CFD) is not routinely used in this area. Both academia and industry anticipate significant advancements in the integration of CFD methodologies by the year 2030. Aiming at conducting shape optimization of airfoils based on a high-fidelity CFD solver with the state-of-the-art technique, a mechanism-driven reinforcement learning framework is proposed in this paper.
@@
The traditional method pioneered by Jameson [5, 6] combined the optimization problem with adjoint equations which has been proven to be very efficient. However, the analysis and application of optimization techniques in the industry can not be generalized to more complicated target functional easily[4]. Combining CFD methods with optimization offers some distinct advantages. For instance, a more precise depiction of flow fields enables a more accurate calculation of target functionals. The advent of techniques such as the dual-weighted residual(DWR)-based mesh adaptation[7, 8, 9] allows for the attainment of more precise target functionals while preserving computational accuracy. Moreover, the enhanced coupling between more accurate geometric representations and PDEs provides superior tools for simulating airfoil geometric deformations through the use of parametric curves and unstructured mesh methods to describe shapes[10, 11]. However, PDEs-constrained optimization also presents various challenges, including the high computational cost associated with solving PDEs[12] with complex geometries, non-linearities inherent in many physical systems, and application issues for ensuring that optimized shapes are physically realizable[13].
@@
With the development of artificial intelligence in recent years, associating the neural networks with traditional computational fluid dynamics problems becomes feasible in shape optimization of airfoil[14, 15]. For example, the Bézier-GAN model in [16, 17] and the conditional variational autoencoder [18] have been adopted to assess the performance of the shape of airfoils. Moreover, the recent advancements in reinforcement methodologies, such as deep reinforcement learning[19], have significantly expanded its applicability and effectiveness in handling high-dimensional state and action spaces.
@@
The distinctive advantages of reinforcement learning, compared with other machine learning algorithms, lie in its ability to make sequential decisions and learn optimal policies directly from interaction with the environment[20]. This is especially pertinent in the realm of control problems governed by partial differential equations[21, 22], where the dynamics are complex and highly nonlinear. Different from traditional machine learning approaches, reinforcement learning thrives in an exploration-based learning environment, making it uniquely suited for optimizing shapes and controls in PDEs-driven systems without the need for extensive pre-processed data. These developments, combined with our group’s work in exploring advanced PDE solvers[23, 24, 25, 10], have laid the groundwork for implementing a PDE solver-based reinforcement learning framework.
@@
Even though reinforcement learning has made significant strides in various applications, its integration into shape optimization, particularly within the context of PDEs-based systems, remains nascent. Ever since in [26], reinforcement learning has been applied to morphing within an airfoil design context. Initially, the application was constrained by the consideration of only four variables for airfoil morphing. With the advent of computational advancements, a direct shape optimization methodology utilizing reinforcement learning was introduced in [27]. Given that numerical simulations in CFD are time-consuming, the number of parameters is still far from the practical requirements. A transfer learning technique [28] is implemented to expedite the learning process. However, such a technique is still not a robust strategy without a mechanism involved. In our research[29, 30, 31, 32], we have integrated the DWR-based mesh refinement strategy, a method that can generate a high quality mesh for the calculation of target functional. It saves time significantly, which supports the subsequent framework conducted with a large number of parameters.
Challenges in developing this framework combining reinforcement learning and PDE solvers are from the following aspects.
• The absence of efficient automation techniques in DWR-based mesh adaptation processes leads to critical challenges, particularly as significant geometric deformations occur. During the iterative process of reinforcement learning, the PDE solver might be invoked thousands of times. Manual intervention at each iteration is impractical[33]. In [34, 32], the automation of DWR-based mesh adaptation is discussed. Further improvements should be implemented to guarantee a highly robust strategy even when the geometry is extraordinarily changed. Otherwise, the accuracy cannot be preserved, which will affect the agent’s learning.
• The deformation conducted within the unstructured mesh will bring potential mesh tangling issues[35]. The change of geometry may be stochastic initially, where the quality of meshes can not be guaranteed. Besides, the curvature may change rapidly along the deformed shape. If the curvature variation occurs within the elements, it is challenging to capture the geometry correctly.
• Tackling the issues posed by controlling geometries with a large number of parameters is challenging[36, 37]. The primary difficulty with a high parameter count is that it complicates the simulation process. Furthermore, while a more precise control over geometry enhances manipulability, it also poses challenges to maintaining a regular global shape.
• Designing a framework that efficiently meets varied optimization goals is inherently complex, requiring the creation of mechanisms specifically crafted for each unique task. The effectiveness of DWR-based mesh adaptation methods is inherently related to the current state of theoretical advancements[7]. The challenge is further compounded when dealing with more complex target functionals, whose resolution poses significant difficulties. Furthermore, translating traditional optimization problems into reinforcement learning frameworks introduces additional complexities. Ensuring the equivalence of primal optimization objectives and the result of reinforcement learning is paramount. This requires the analysis of the reward function[38], necessitating it to be closely compatible with the underlying optimization context.
In this paper, we propose a novel mechanism-driven reinforcement learning, aiming at enhancing the shape optimization of airfoils. This approach effectively manages high-dimensional geometric parameters over unstructured meshes for different objectives, addressing several key challenges in the field. Prior to this work, there has been limited work combining the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm with traditional PDE solvers for direct airfoil design optimization in such high-dimensional spaces. It is worth mentioning that we have analyzed the equivalence of the learning process and objectives based on specific reward function.
We integrate an automatic algorithm leveraging the DWR-based mesh adaptation method, facilitating automation essential for iterative optimization and ensuring time efficiency without compromising accuracy. Subsequently, we mitigate potential mesh tangling during airfoil deformation by incorporating Laplacian smoothing, which prevents irregular triangle accumulation and reduces finite element analysis errors due to mesh irregularities. Then, the geometric manipulation utilizes up to 132 sample points for precise actions, coupled with a regularization strategy for control points generation. This approach ensures regular shape transitions and captures intricate curvature details, enhancing geometrical fidelity through Bézier fitting for effective dimensionality reduction. Lastly, our framework employs the enhanced TD3 algorithm, tailored for high-dimensional control problems. It supports the optimization task with a well-aligned reward function, promoting behaviors yielding higher rewards and penalizing undesirable actions, thereby streamlining the learning process towards achieving optimization goals efficiently.
@@
The rest of this paper is organized as follows. In section 2, we briefly introduced the PDEs-constrained optimization, fundamental notations and the architecture of the automatic DWR-based mesh adaptation method. In section 3, a basic process of the reinforcement learning based on the shape optimization of airfoil is introduced. Additionally, issues for setting up the learning process which originate from the high dimensionality and unstructured mesh are addressed. In section 4, the algorithm of Twin Delayed DDPG is constructed. Further details of each component of the algorithm, analysis of the equivalence between the learning process and optimization task are elaborated. In section 5, we present the numerical results. The future improvements are discussed in section 6.
2 PDEs-constrained optimization
This study focuses on the shape optimization of airfoils using PDEs-constrained optimization techniques. Consider an objective function $J(\mathbf{u})$ that we seek to minimize, subject to a constraint imposed by a governing equation $R(\mathbf{u})=0$. The optimization problem can be mathematically formulated as follows:
$\min_{\mathbf{u}}J(\mathbf{u})\quad\text{subject to}\quad R(\mathbf{u})=0.$ (1)
In this work, we integrate PDEs-constrained optimization within a reinforcement learning framework. Our approach involves navigating the complex landscape of PDEs constraints to accurately achieve the desired target functional. We utilize the DWR method for mesh adaptation, which will be elaborated upon in this section. This method enhances our ability to refine the mesh dynamically based on error estimates that contribute directly towards improving the accuracy of the objective function $J$.
2.1 Steady Euler equations
@@
In our previous works [30, 29, 24], we have developed a solver capable of handling the steady Euler equations. This solver not only facilitates the computation of these equations but also ensures the accurate resolution of the target functional, laying a solid foundation for our shape optimization endeavors. A brief introduction will be given in this section.
@@
The steady Euler equations help us understand inviscid flow behavior and are capable of addressing the shape optimal design problem[39]. These equations are represented in a conservative form for the inviscid two-dimensional steady state,
@@
For instance, we consider the optimization of airfoil shapes with the primary objective of drag minimization. We define the drag as a specific quantity of interest $J_{d}(\mathbf{u})$, which is to be minimized. The mathematical formulation of the optimization problem is given by:
$\min_{\mathcal{S}}J_{d}(\mathbf{u})\quad\text{subject to}\quad\nabla\cdot\mathcal{F}(\mathbf{u})=0,\quad\text{in}\ \Omega_{\mathcal{S}},$ (5)
where $\mathcal{S}$ denotes the shape of the airfoil, and $\Omega_{\mathcal{S}}$ represents the fluid domain bounded by the airfoil shape $\mathcal{S}$. The governing equation $\nabla\cdot\mathcal{F}(\mathbf{u})=0$ encapsulates the flow dynamics within the domain, ensuring that the flow is modeled accurately according to the physical constraints.
Solving the PDEs and obtaining a precise target functional is our first concern. Then we discretize the domain $\Omega_{\mathcal{S}}$. Consider $\Omega_{\mathcal{S}}$ as a domain in $\mathbb{R}^{2}$ with boundary denoted by $\Gamma_{\mathcal{S}}$. We divide $\Omega_{\mathcal{S}}$ into distinct control volumes, $K_{i}$, via a shape-regular subdivision. Then the mesh under the current mesh size is denoted as $\mathcal{K}_{h}$. Each control volume is referred as $K_{i}$. The intersection between any two elements $K_{i}$ and $K_{j}$, is labeled as $e_{i,j}$, satisfying $e_{i,j}=\partial K_{i}\cap\partial K_{j}$. The edge $e_{i,j}$ relative to $K_{i}$ is equipped with a unit outer normal vector, $n_{i,j}$. This setup allows us to redefine the weak form of Euler equations through the divergence theorem as:
@@
To tackle the nonlinear Equation (7), we employ the Newton-GMG framework. Details can refer to [40, 23, 25]. To obtain a precise value of the target functional, we turn to use the DWR-based mesh adaptation as part of the PDE-solver.
2.2 An automatic DWR-based adaptation for the target functional calculation
In [29, 31], we elaborated a framework innovated by Darmofal [41, 42]. To incorporate this technique within our framework, we present an overview of the methodology. Denote solutions on the coarse mesh as $\mathbf{u}_{H}$ and the finer mesh as $\mathbf{u}_{h}$. The initial goal is computing the integration of $J(\mathbf{u})$ over a fine mesh, denoted as $J_{h}(\mathbf{u}_{h})$. A multi-variable Taylor series expansion method is adopted as follows: